OpenStudy (anonymous):

Which quadrant do they belong in? 22. tan theta is negative 23. sec theta is negative 24. sin theta is negative 25. cos theta is negative 26. sin theta > 0 and cos theta <0 27. tan theta > 0 and sec theta < 0 28. csc theta < 0 and cos theta > 0 29. cot theta > 0 and sin theta > 0

OpenStudy (campbell_st):

the diagram shows what ratios are positive in each quadrant.  use this to answer your questions.... remember sec = 1/cos csc = 1/sin cot = 1/tan so the rules still apply